The current I is an electrical conductor varies inversely as the resistance R of the conductor. The current is 1/7 ampere when the resistance is 21609 ohms. What is the current when the resistance is 3969 Ohms?
K=1/7*21609=3087
I=3087/R=
How do I use the equation I=3087/R= to find the current when the resistance is 3969 Ohms? I figured out the equation however, I don’t know what to do next.

Answer by rfer(16322)

I=3087/3969
I=0.7777777777 amps
check
E=IR
E=0.14286*21609
E=3087 volts
-----------
E=3969*0.7777777777
E=3087 volts